From GlueXWiki
Jump to: navigation, search

June 23, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Second and third cell status
    • Humidity requirements (related to possible changes of the AC settings)
    • Other: using radioactive source at Blue Crab
  2. Engineering (Bill)
    • Wire deadening Sasha's doc (page 13, Fig.14)
    • Package cabling
    • Other
  3. Electronics update (Fernando, Chris)
  4. Chamber testing at EEL126 (Beni, Lubomir)
    • First cell status
    • DAQ and software
  5. Other


Participants: Eugene, Bill, Tim, Dave, Chris, Mark, Beni, Simon, Casey, and Lubomir.


- Dave (see the Construction Tracking link above): cell#2 is ready, including one end window. Anatoly cleaned the wire frame manually with alcohol. Second end window will be ready tomorrow. Will start very soon stacking the package, but need some parts. Cell#3: position measurements on the wire frame (Caleb and Al made a code that does it automatically, but has not been tested yet), making flaps on one of the cathodes and gluing rigid-flexes on the other one. Cell#4: wire frame is ready for stringing, one of the cathodes is being tensioned and the foil for the second is in preparation.

- We discussed if we can relax the clean room humidity requirements to save on the energy bill. The main concerns are the gluing operations and the Rohacell material being susceptible to elevated humidity. Tim explained: there's no humidity control and therefore it depends on the heat in the room (that's why it was elevated over the last weekend); we could increase the setting temperature from 69 to 70deg F to see the effect. Eugene: first we have to establish the production and then to investigate the effect of the humidity, even 50% is very high. Conclusion: at least for this month we don't want to change the settings.

- Bert Manzlak is trying to help us establishing a Radiation Control Area at Blue Crab and organized discussions with Keith Welch. RadCon are reluctant and are looking into the administrative papers if that is possible at all. According to Bert, a RCA was made off-site long time ago.


- Bill: we have most of the tooling needed for the wire deadening but have to know the radii of the dead areas. Lubomir: originally it was decided to be able to register pions with 1GeV at 1deg which requires upper limits: 2.3cm 3.2cm 3.9 4.6cm (package#1-#4). Alex simulated the rates (doc linked above) and requiring to have at most 130kHz/strip gives lower limits: 2.9, 3.0, 3.0, 3.5cm. There's a conflict for the first package but it can be 2.9cm without affecting the acceptance noticeably resulting in: 2.9, 3.2, 3.9, 4.6cm. Finally, to be able to interchange modules we decided to have 3.0cm, 3.0cm, 3.9cm, 3.9cm radii of the dead areas.

- FDC cabling scheme was discussed on a meeting this week and it was decided to build a mock-up. Bill gave us some details of the cable supports. We discussed the conduit material and if we need spare cables installed. Beni made a point that fixing connectors on the signal cables may require cutting the cables by 1ft to be in the flat part of the cable, as it happened already. That's why we need spare signal cables installed. Eugene regarding the number of the spare cables: we need to cover a sector for each package, say 4 or 8 per package; on the other hand we will have 42 conduits so it make sense to have 42 or 21 spare cables, something to be investigated with the mock-up. As for the HV/LV cables there are ways to fix/split the cables and such spares installed are not needed.

- Bill: the gusset rings might be ready sooner than expected, at the end of the next week.


- Chris: First 75(CDC)+10(FDC) signal cables were tested, expect to have 75 cables each week. Talked to the rigid-flex stuffing company and ordered the rest of the rigid-flexes to be tinned.

Chamber testing

- Lubomir: gas and HV problems with the first cell at 126. On Tuesday, after installing the chamber it started bubbling at 50ccpm total, 90/10 Ar/CO2. The flow was increased to 100ccpm (because the flow controllers are not so precise at that low rates) and the chamber stopped bubbling. It turned out the differential pressure inside was 45-47Pa, indicating blockage of the exhaust pipe. After reducing the pressure down to 30Pa it started bubbling again. At the moment the chamber is bubbling at 75ccpm and 40Pa. Bill, Mark and Casey discussed the current and previous procedures for making the gas spacer and possible problems. Bill has assumed a pressure of 30Pa, he is afraid the end window mylar can peel off. Bill will investigate possible solutions to keep the pressure bellow some level. HV problems: even we were able to operate the chamber at 2200V (except one sector at 2100V) at Blue Crab, now at 126 we could have at most 1700V. It could be temperature/humidity problem or something else; the conditioning will continue.

- Beni had problems with the DAQ, maybe caused by overheating the crate. We discussed options for installing additional fans.